ELECTRONIC SPECTROSCOPY OF THE A^{1}A′′-X^{1}A′ SYSTEM OF CHCl AND CDCl

\underline{C. TAO}, M. DESELNICU, C. MUKARAKATE AND S. A. REID, Department of Chemistry, Marquette University, Milwaukee, WI 53201-1881.

We report a comprehensive new study of the electronic spectroscopy of the A^{1}A′′-X^{1}A′ system of CH^{35}Cl, CH^{37}Cl, CD^{35}Cl and CD^{37}Cl in the region 450-750 nm. For each isotopomer, cold bands in the progressions 2_{0}^{n}, 2_{0}^{n}3_{0}^{m} (m=1-2), 1_{0}^{1}2_{0}^{n}, and 1_{0}^{1}2_{0}^{n}3_{0}^{m} (m=1-2) were observed. The spectra were measured under jet-cooled conditions using a pulsed discharge source, and rotationally analyzed to yield precise values for the band origins and rotational constants. The derived A^{1}A′′ state vibrational intervals are combined in Dixon plots with previous experimental results^{a}^{,}^{b}^{,}^{c}^{,}^{d} to derive barriers to linearity for the 2^{n} progression. The A^{1}A′′ C-H (C-D) stretching frequency was determined for the first time, and the observed ν_{2} and ν_{3} dependence of the CH(D)^{35}Cl-CH(D)^{37}Cl isotope splitting in the A^{1}A′′ state is investigated.

aA.-J. Merer and D. N. Travis, Can. J. Phys. 44, 525 (1966).

bB.-C. Chang and T. J. Sears, J. Mol. Spectrosc. 173, 391 (1995).

cB.-C. Chang and T. J. Sears, J. Chem. Phys. 102, 6347 (1995).

dA. Lin, et al., J. Mol. Spectrosc. 214, 216 (2002).